OpenStudy (anonymous):

Antibiotic B, is taken every 6 hrs and has an effectiveness factor of 90%. Determine the bacterial population after three days (72h) in a patient who has an intial bacterial population of 2000 and who is being treated with antibiotic B

OpenStudy (matt101):

This is basically a decay problem. Every 6 h, the bacterial population decreases to 10% its starting value. The general equation would be: p(t) = p(0.1)^(t/k) Where p(t) is the population at time t, p is the starting population, t is the total time elapsed, and k is the length of each halving cycle (if this were an atomic decay question, that would mean the half-life). Plug in your numbers and see what you get.

OpenStudy (anonymous):

y=2000(.1)^t t=72/6 t=12 y=2000(.1)t^12 y= 2x10^-9 Hmm doesnt seem right, but Matt will probably have it right :)

OpenStudy (matt101):

Yeah you're right...that seems sorta off. Is there a specific definition for "effectiveness factor" that's different from what we used?

OpenStudy (matt101):

Also @Brownguy - thanks for the vote of confidence lol

OpenStudy (anonymous):

Matt would it be an inconvience if you solve it Completely cause i am still stuck?

OpenStudy (anonymous):

@matt101